Find the height of a triangle with a base of 30m and an area of 45m

Answer by rothauserc(4718)

area(A) of triangle = (1/2)*b*h where b is the base and h is the height
we are given,
45 = (1/2) * 30 * h
multiply both side of = by 2
90 = 30*h
divide both sides of = by 30
h = 3
the height of our triangle is 3 m
check the answer
45 = (1/2) * 30 * 3
45 = (1/2) * 90
45 = 45
our answer checks